30. “Double taxation refers to the LR.S. first taxing a corporations revenue and then its profits. 31. All else the same, t True or false
Answer #1

30.False.Double taxation refers to taxing the same income twice in different countries

31.True

32.False as It is not clearly mentioned that shares are sold through an IPO.If it is sold through IPO(Intial Public offering),than it is a primary market transaction otherwise it an secondary market transaction.

33.True.as it is sold through stock market it is an secondary market transaction

35.True.Debt equity ratio = debt/equity=1/4=0.25:1

Equity =5-1=4 considered